You intend to draw a random sample of size n=526n=526 from a population whose parameter is p = 0.53

What is the standard deviation of the distribution of sample means?
(Report answer accurate to 2 decimal places.)
Answer #1

Solution

Given that,

p = 0.53

1 - p = 1-0.53=0.47

n = 526

mean = = p =0.53

standard deviation = =  [p ( 1 - p ) / n] =   [(0.53*0.47) /526 ] = 0.02